OpenStudy (anonymous):

Using the elimination method how do i solve . 2x+3y=13 x-3y= -7

OpenStudy (kropot72):

First add the equations to eliminate the terms in y. Please do this and post the result.

OpenStudy (anonymous):

3x=6 x=2 2-3y=-7 -3y=-9 y=3

OpenStudy (kropot72):

Sorry we couldn't work thru to the solution Seabee :(

OpenStudy (anonymous):

no , thank you actually im trying to work on it myself now

OpenStudy (anonymous):

that's good :D

OpenStudy (kropot72):

Very glad to hear it Seabee :)